OpenStudy (anonymous):

Your Cousin is 26 yr older than your sister. The product of their ages is 560 . How old is your sister?

hero (hero):

(x)(x+26) = 560

OpenStudy (anonymous):

Show work haha?

hero (hero):

x^2 + 26x - 560 = 0

hero (hero):

Quad Equation

OpenStudy (anonymous):

so it is 560 or 0?

OpenStudy (anonymous):

or you could just ask , lol

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

ask them their age

OpenStudy (anonymous):

its a math problem?

OpenStudy (anonymous):

what????

hero (hero):

40 and 14, sorry

OpenStudy (anonymous):

x^2+26x-560=(x-40)(x+14) x=40 or -14 -14 is not an age so ur sister is 40

hero (hero):

beat you again jah

hero (hero):

actually, the sister is 14

OpenStudy (anonymous):

yup read it backwards lol

OpenStudy (anonymous):

WHATS THE ANSWER , SHOW WORK hahahaha ??

hero (hero):

If you don't know quad equations, can't help you

OpenStudy (anonymous):

x=sisters age x+26=cousins age (x)(x+26)=560 x^2+26x=560 x^2+26x-560=0 solve for x gives ur sis age (x+40)(x-14)=0 x=14 or -40 since -40 is not an age ur sis is 14 now find cousins age x+26 x=14 so 14+26=40 ur cuz is 40